# Exploring Large-Volume GAGG Scintillators for use in MeV Gamma-Ray Astrophysics

### v3

@@ Abstract
Gamma-ray astrophysics in the MeV band is an exciting field in astronomy due to its potential for multi-messenger astrophysics. It has, however, remained under-explored when compared to other wavelengths. One reason for this observational gap is the difficulties with measuring these high-energy photons and the requirement of large amounts of detection material. In this work, we investigate the usage of large-volume GAGG scintillators for use as a calorimeter in future MeV telescopes. We developed a $5\times 5$ array calorimeter utilizing $1\times 1\times 6\ \mathrm{cm}^{3}$ GAGG crystals with C-series SiPM readout. We tested the calorimeter at the High Intensity Gamma-ray Facility (HIGS) with monoenergetic beams ranging from $2-25\ \mathrm{MeV}$. Finally, we also investigated larger $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals and characterize their response across their depth when their surface treatment are either polished or frosted.
@@ 1 Introduction
The gamma-ray astrophysical sky in the MeV range remains largely under-explored. The most recent large-scale mission dedicated to measuring and imaging the MeV gamma-ray sky was the Compton Gamma-Ray Observatory (CGRO), which was de-orbited in 2000 [1]. Since then, the community has proposed several Compton and/or pair-production instruments to address what is known as the ‘MeV gap’. Most pair-production instruments are based on a two-subsystem design: a tracker to trace out the movement of the cascade of particles from the gamma-ray interaction and a calorimeter to capture the ensuing electromagnetic showers. This tracker-calorimeter design is utilized by current missions such as the Fermi Space Telescope-LAT [2] and AGILE [3]. Missions that are in the concept and development phase include AMEGO-X [4], ASTROGAM [5], SMILE-2+ [6], VLAST [7], and GECCO [8].
This work investigates the use of gadolinium aluminum gallium garnet (GAGG) based scintillator to act as a high-energy calorimeter for potential usage in gamma-ray telescopes. GAGG presents several advantages over CsI, as outlined in Tab. 1. First, it has superior energy resolution. Next, it is significantly denser than most other scintillators, which results in a more compact design. A faster decay time also allows for operation at higher count rates.
@@ 2 Development of a $5\times 5$ array GAGG Calorimeter
In this work, we utilize high light-yield GAGG from Kinheng. The prototype in this section uses $1\times 1\times 6\ \mathrm{cm}^{3}$ crystals, each one wrap with a diffuse white reflector (Tetratex) [21], and place a $6\times 6\ \mathrm{mm}^{2}$ C-series SiPM by onsemi [22] on each end of the crystal. Fig. 1 shows the bare crystal and a fully developed crystal element.
@@ 2.1 Gamma-ray response
where $Q_{top}$ and $Q_{bottom}$ are the measured signals from the top and bottom SiPM on the same crystal.
@@
The measured resolution at 662 keV is around 8%, which is poorer than the general advertised value. This is most likely due to the difference in geometry as this GAGG is longer than mainstream crystals that are on the market.
@@ 3 High-Energy Gamma-Ray Response of the Calorimeter
Fig. 7 shows the crystal-by-crystal response to different irradiation configurations relative to the 5 MeV beam. Fig. 7(a) plots the front irradiation while Fig. 7(b) plots side irradiation. The two irradiations show the expected full energy and escape peaks in the targeted crystals, with the surrounding crystals showing significant continuum and 511 keV depositions from the ensuing cascades. Also detected is the 1.46 $\mathrm{MeV}$ from 40K present in the background. When comparing to the summed spectra in Fig. 6, we observe a very different peak-to-escape ratio with the double escape peak being the most prominent feature.
@@
Figure 7: Crystal-to-crystal response to a 5 MeV beam while the calorimeter is oriented at different poses.
Fig. 8 plots all the acquired spectra from the HIGS campaign. Full energy peaks are clearly visible for the 2, 5, and 7 $\mathrm{MeV}$ beams. Also featured are the single and double escape peaks. However, runs with 15 and 25 $\mathrm{MeV}$ do not present any evidence of escape peaks, likely due to the low full-energy efficiency and full-containment of the ensuing spallation. The experimental response matches the simulated one, presented in A.
@@ 4 Depth of Interaction response study of GAGG Crystals
We expand our investigation of GAGG by exploring longer crystals that are $1\times 1\times 8\ \mathrm{cm}^{3}$ in volume. Moving to the longer crystals is of interest as it increases the detection efficiency and full-energy containment. In Fermi’s LAT calorimeter, all the sides were roughened to introduce a depth-dependent light output when read out on each end of the crystal. In this section, we investigate the difference in depth of interaction response when the sides of the crystals are either polished or frosted. Fig. 9 shows the roughened crystal (on the right), next to a polished version (left). Each crystal is packaged and readout like the crystals described in Sec. 2.
@@ 4.1 Collimated Depth Measurements
In this section, we conducted collimated measurements at each depth to verify the scintillation properties across the entire crystal and the depth reconstruction scheme. A fan beam is produced by using two lead bricks separated by an American penny with a ${}^{137}\mathrm{Cs}$ source placed on the other end. The assembly is mounted on a robotic gantry that is programmed to move in $5\ \mathrm{mm}$ steps and scan along the length of the GAGG crystal. Fig. 10 illustrates the experimental setup.
@@
Like in the Fermi LAT Calorimeter, we calculate the depth of interaction (DOI) using Eq. 2.
@@
We can then make a DOI calculation as a function of the collimator position and plot it in Fig. 12. The figure also plots the calculated DOI for a polished crystal. The frosted crystal displays a very linear response to the different depths indicating that applied surface roughening is sufficient. On the other hand, the polished crystal displays a negligible amount of depth response. Nevertheless, both crystals display a fully active scintillator.
@@ 4.2 Flood Irradiation Response
The energy and depth are then calibrated using techniques described in [24]. Fig. 14 plots the corrected response, highlighting the ${}^{137}\mathrm{Cs}$ peak, where the top and bottom rows plot the frosted and polished crystals respectively.
Figure 14: Corrected Depth Energy response of the $1\times 1\times 8\ \mathrm{cm}^{2}$ GAGG crystals. The top row plots the 3 different frosted crystals while the bottom row plots 3 polished crystals.
@@ 5 Conclusion and Future Work
This work explores the usage of GAGG as a high-energy calorimeter for a MeV range gamma-ray telescope. We built a $5\times 5$ array of GAGG crystals, each $1\times 1\times 5\ \mathrm{cm}^{3}$ in size. This system was then characterized with monoenergetic gamma rays ranging from $2$ to $25\ \mathrm{MeV}$. The results show good agreement with simulations. Next, we investigate the depth response of $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals that are either frosted or polished on all sides. Collimated measurements show that the longer $8\ \mathrm{cm}$ crystals are fully active and display a near-linear response when using standard DOI calculations.
We are currently developing a larger $8\times 8$ array calorimeter using the $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals. In addition, we are developing SiPM readouts that field two different SiPM species such that their different gains will increase the dynamic energy range of the readout [26]. The ‘dual-gain’ SiPMs are currently under fabrication by Fondazione Bruno Kessler [27].
@@ Appendix A Simulated Response of the GAGG Calorimeter
We modeled the $5\times 5$ GAGG calorimeter using the SWORD simulation package [28]. In addition to the calorimeter (shown in Fig. 2(a)), the ComPair system was also modeled to simulate the HIGS beam test conditions. The plots in this section show the GAGG calorimeter’s response to $2,5,7,15$ and $25\ \mathrm{MeV}$ monoenergetic gamma rays. The simulated energy blurring is accomplished using the mean energy resolution presented in Fig. 4. However, we did not account for the crystal to crystal variations in energy response.
@@
For energies ranging from $2-7\ \mathrm{MeV}$, a clear full energy peak is visible along with the two escape peaks, which mirrors what is observed during experimentation. The $15\ \mathrm{MeV}$ simulated run shows very modest full energy and escape peaks. On the other hand, the experimental result shows a skewed left continuum with no obvious features. This is likely a result of the simulation not accounting for the crystal to crystal variations. The $25\ \mathrm{MeV}$ run shows a continuum with no spectral peaks, similar to the experimental run.

### v4

@@ Abstract
Gamma-ray astrophysics in the MeV band is an exciting field in astronomy due to its potential for multi-messenger astrophysics. It has, however, remained under-explored when compared to other wavelengths. One reason for this observational gap is the difficulties with measuring these high-energy photons and the requirement of large amounts of detection material. In this work, we investigate the usage of large-volume GAGG scintillators for use as a calorimeter in future MeV telescopes. We developed a $5\times 5$ array calorimeter utilizing $1\times 1\times 6\ \mathrm{cm}^{3}$ GAGG crystals with onsemi C-series SiPM readout. We tested the calorimeter at the High Intensity Gamma-ray Facility (HIGS) with monoenergetic beams ranging from $2-25\ \mathrm{MeV}$. Finally, we also investigate larger $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals and characterize their response across their depth when their surface treatment is either polished or frosted.
@@ 1 Introduction
The gamma-ray astrophysical sky in the MeV range remains largely under-explored. The most recent large-scale mission dedicated to measuring and imaging the MeV gamma-ray sky was the Compton Gamma-Ray Observatory (CGRO), which was de-orbited in 2000 [1]. Since then, the community has proposed several Compton and/or pair-production instruments to address what is known as the ‘MeV gap’. Most pair-production instruments are based on a two-subsystem design: a tracker to trace the movement of the cascade of particles from the gamma-ray interaction and a calorimeter to capture the ensuing electromagnetic showers. This tracker-calorimeter design is utilized by current missions such as the Fermi Space Telescope-LAT [2] and AGILE [3]. Missions that are in the concept and development phase include AMEGO-X [4], ASTROGAM [5], SMILE-2+ [6], VLAST [7], and GECCO [8].
This work investigates using gadolinium aluminum gallium garnet (GAGG) based scintillator to act as a high-energy calorimeter for potential usage in gamma-ray telescopes. GAGG presents several advantages over CsI, as outlined in Table 1. First, it has superior energy resolution. Next, it is significantly denser than most other scintillators, which results in a smaller instrument without compromising the high-energy response. A faster decay time also allows for operation at higher count rates. Moreover, some reports declare a lower light-yield degradation due to radiation damage [9]. The table also reveals several disadvantages. Perhaps the most significant downside is its afterglow properties in that it is larger than most scintillators and endures for longer [10, 9].
@@ 2 Development of a $5\times 5$ array GAGG Calorimeter
In this work, we utilize high light-yield GAGG:Ce from Kinheng. The prototype in this section uses $1\times 1\times 6\ \mathrm{cm}^{3}$ crystals. Each crystal is wrapped twice with a $216\ \mu\mathrm{m}$ thick diffuse white reflector (Tetratex) [23]. It was chosen based on our previous experience and its overall favorable reflective properties [22, 24]. We placed a $6\times 6\ \mathrm{mm}^{2}$ C-series SiPM by onsemi [25] on each end of the crystal. The C-series SiPM was also chosen due to our familiarity with it from the NeRDI-1A instrument [26]. Fig. 1 shows the bare crystal (top) and a fully built crystal element (bottom).
@@ 2.1 Gamma-ray response
where $Q_{top}$ and $Q_{bottom}$ are the measured signals from the top and bottom SiPM on the same crystal. Although different energy treatments exist, we utilize this form based on our experience from the ComPair CsI balloon instrument [22].
@@
The measured resolution at 662 keV is around 8%, which is poorer than the general advertised value. This is most likely due to the difference in geometry as this GAGG is longer than mainstream crystals on the market. Moreover, GAGG is known to be significantly more self absorbing when compared to other scintillators [29]. The degradation of scintillator resolution with geometry has been extensively documented [30, 31, 32].
@@ 3 High-Energy Gamma-Ray Response of the Calorimeter
Fig. 7 shows the crystal-by-crystal response to different irradiation configurations relative to the 5 MeV beam. Fig. 7(a) plots the front irradiation while Fig. 7(b) plots side irradiation. The two irradiations show the expected full energy and escape peaks in the target crystals. The double escape peak is enhanced when viewing the spectrum from a single crystal compared to the summed response in Fig. 6. More specifically, the double escape peak is the most prominent feature in the single crystal spectra. The summed response has the single escape peak as the most prominent with the full-energy peak being a close 2nd. The surrounding crystals contain significant continuum and 511 keV depositions from the ensuing cascades. The high energy response in the surrounding crystals most likely originates from the off-axis artifacts of the beam.
@@
Figure 7: Crystal-to-crystal response to a 5 MeV beam while the calorimeter is oriented at different poses. The y-axes of the subplots are not shared with the range left floating. The reader can infer the number of counts in each subplot from the statistical noise in each trace.
Fig. 8 plots all the acquired spectra from the HIGS campaign. Full energy peaks are clearly visible for the 2, 5, and 7 $\mathrm{MeV}$ beams. Also featured are the single and double escape peaks. However, runs with 15 and 25 $\mathrm{MeV}$ do not present any evidence of escape peaks, likely due to the low full-energy efficiency and low full containment of the ensuing spallation. The experimental response matches the simulated one, presented in A.
@@ 4 Depth of Interaction response study of GAGG Crystals
We expand our investigation of GAGG by exploring longer crystals that are $1\times 1\times 8\ \mathrm{cm}^{3}$ in volume. Moving to the longer crystals is of interest as it increases the detection efficiency and full-energy containment. In Fermi’s LAT calorimeter, two of the four sides were roughened to introduce a depth-dependent light output when read out on each end of the crystal [34]. In this section, we investigate the difference in depth of interaction response when the sides of the crystals are either polished or frosted. Fig. 9 shows the roughened crystal (on the right), next to a polished version (left). Each crystal is packaged and read out like the crystals described in Sec. 2.
@@ 4.1 Collimated Depth Measurements
In this section, we conducted collimated measurements at each depth to verify the scintillation properties across the entire crystal and the depth reconstruction scheme. A fan beam is produced by using two lead bricks separated by a $1.52\ \mathrm{mm}$ thick American penny with a ${}^{137}\mathrm{Cs}$ source placed on the other end. The assembly is mounted on a robotic gantry that is programmed to move in $5\ \mathrm{mm}$ steps and scan along the length of the GAGG crystal. Fig. 10 illustrates the experimental setup.
@@
We calculate the depth of interaction (DOI) using Eq. 2.
@@
We can then make a DOI calculation as a function of the collimator position and plot it in Fig. 12. The figure also plots the calculated DOI for a polished crystal. The frosted crystal displays a very linear response to the different depths indicating that the applied surface roughening is sufficient. On the other hand, the polished crystal displays a negligible amount of depth response. Nevertheless, both crystals display a fully active volume in that their entire length was sensitive to gamma rays.
@@ 4.2 Flood Irradiation Response
The energy and depth are then calibrated using techniques described in [28]. Fig. 14 plots the corrected energy-depth response, highlighting the ${}^{137}\mathrm{Cs}$ peak, where the top and bottom rows plot the frosted and polished crystals respectively.
Figure 14: Corrected depth-energy response of the $1\times 1\times 8\ \mathrm{cm}^{2}$ GAGG crystals. The top row plots the 3 different frosted crystals while the bottom row plots 3 polished crystals.
We plot the spectra of the crystals in Fig. 15 to highlight their gamma-ray response. We show that the depth response of the frosted crystals is removed. Due to the crystal geometry, the full-energy peak-to-continuum ratio is rather poor. Table 2 summarizes the resolution of each crystal. We show that the resolution is slightly degraded when compared to the $6\ \mathrm{cm}$ crystal. A possible explanation of why the frosted crystals outperform the polished ones is that the depth corrections can correct any regional defects in the crystal. However, a larger sample of crystals is required to declare a conclusive statement.
Figure 15: Energy response of the $1\times 1\times 8\ \mathrm{cm}^{2}$ GAGG crystals using the calibration matrix. The top row plots the 3 different frosted crystals while the bottom row plots 3 polished crystals.
Table 2: Percent full width at half maximum at $662\ \mathrm{keV}$ displaying the resolution of the $8\ \mathrm{cm}$ long crystals.
Best Crystal [%] Average [%]
Frosted 8.2 8.5 $\pm$ 0.2
Polished 8.5 8.8 $\pm$ 0.3
@@ 5 Conclusion and Future Work
This work explores the usage of GAGG as a high-energy calorimeter for a MeV range gamma-ray telescope. We built a $5\times 5$ array of GAGG crystals, each $1\times 1\times 6\ \mathrm{cm}^{3}$ in size. This system was then characterized with monoenergetic gamma rays ranging from $2$ to $25\ \mathrm{MeV}$. The results show good agreement with simulations and present an initial proof of concept.
Next, we investigate the depth response of $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals that are either frosted or polished on all sides. Collimated measurements show that the longer $8\ \mathrm{cm}$ crystals are fully active showing that the long GAGG crystals are sensitive to gamma rays throughout the entire volume. Moreover, polishing the crystals on all 4 sides resulted in the desirable depth-dependent signals when the fully polished crystal yielded negligible depth response. The polished crystals displayed a near-linear response depth response when using Eq. 2 to calculate the depth, reducing the need for advanced calibration and non-linearity corrections. Future iterations of this finger calorimeter design would opt for crystals where the 4 long faces are polished. The depth response is desired when attempting to reconstruct the particle tracks or ensuring electromagnetic showers for the targeted gamma-ray telescope applications.
We are currently developing a larger $8\times 8$ array calorimeter using the $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals. In addition, we are developing SiPM readouts that field two different SiPM species such that their different gains will the dynamic energy range of the readout [35]. The ‘dual-gain’ SiPMs are currently under fabrication by Fondazione Bruno Kessler [36].
@@ Appendix A Simulated Response of the GAGG Calorimeter
We modeled the $5\times 5$ GAGG calorimeter using the SWORD simulation package [37]. In addition to the calorimeter (shown in Fig. 2(a)), the ComPair system was also modeled to simulate the HIGS beam test conditions. The plots in this section show the GAGG calorimeter’s response to $2,5,7,15$ and $25\ \mathrm{MeV}$ monoenergetic gamma rays. The simulated energy blurring is accomplished using the mean energy resolution presented in Fig. 4. However, we did not account for the crystal-to-crystal variations in energy response.
@@
For energies ranging from $2-7\ \mathrm{MeV}$, a clear full energy peak is visible along with the two escape peaks, which mirrors what is observed during experimentation. The $15\ \mathrm{MeV}$ simulated run shows very modest full energy and escape peaks. On the other hand, the experimental result shows a skewed left continuum with no obvious features. This is likely a result of the simulation not accounting for the crystal-to-crystal variations. The $25\ \mathrm{MeV}$ run shows a continuum with no spectral peaks, similar to the experimental run.